OpenStudy (anonymous):

Evaluate the expression without using a calculator: sin pi/6 + cos pi/6

OpenStudy (campbell_st):

use the exact ratio triangle  then \[\sin(\frac{\pi}{6}) = \frac{1}{2} \] \[\cos(\frac{\pi}{6}) = \frac{\sqrt{3}}{2} \] so the solution is \[\frac{1}{2} + \frac {\sqrt{3}}{2} = \frac{?}{2}\] you just need to finish it off

OpenStudy (anonymous):

okay! thanks